Work is our business (WOB) has a shop with three (3) work centers (WC). Each of the three WCs are available to work 8 hours a day 5 days a week. WOB worked 20 days last month. WOB does a variety of jobs, but last month they did a special order for their best customer and that was the only work they did. Assume that WOB works Monday through Friday and that the truck arrives to take shipments after the WCs finish processing for the day, so that all the work on Friday is available to ship on Friday. Last month they produced only a product called PXXX for this customer. PXXX required processing time (T1) at WC! of 2.2 hours for a single PXXX. After leaving WC1, job PXXX went to WC2 where it's processing time (T2) was about 1 hours. Job PXXX then went to WC3 were it's processint time (T3) was 1.5 hours. When job PXXX completed processing at WC3 it returned to WC2 for one final pass, where its processing time was again, 1 hours. When it finished the second time on WC2 it went to shipping. Shipping accumulated the completed PXXX into pallets and prepared to ship on Friday afternoon. WOB has 2 machines at WC1 and it has 2 machines at WC@ and it has 3 machines at WC3.

1. The total processing time (Tprocess) for each individual product PXXX from its start at WC1 through completion on WC2 (i.e., after the second time at WC2) was:

a. 2 hours

b. 3 hours

c. 4.5 hours

d. 5 hours

e. 5.5 hours

f. 6 hours

g. 8 hours

2. If each work center has 100% uptime on all of its machines, what would be the maximum number of PXXX that can be produced by WOB last month?

3. if 20 jobs were started and completed last week by WOB, what was the utilization of the machines at Work Center 2?

4. Which workstation is the bottleneck when producing 20 jobs a week?
Answer #1

Answer 1. Total processing time Tprocess = T1+T2+T3+T4 = 2.2+1+1.5+1 = 5.7 hours. None of the options provided here match the answer, however the concept is to add all the time required to process one product.

Answer 2. In the above process, at 100% utilization, the bottleneck is at WC1 because it is taking the maximum time to finish the process. This means that there will be one product coming out of this processing line every 2.2 hours.

Therefore, to calculate the maximum number of PXXX produced last month:

No. of days (D) = 20 days

Hours per day (H) = 8 hours

Hours worked last month = D x H = 20 x 8 = 160 Hours

Total PXXX produced last month = 160 / 2.2 = 72.7 units

Answer 3. Days per week = 5

Jobs done last week = 20

Hours of work done at WC2 = 20 x (1+1) = 40 hours

Available hours at WC2 last week = 5 days x 8 hours per day = 40 hours

Utilization of machines at WC2 = (40/40) x 100 = 100%

Answer 4. When producing 20 jobs a week, the bottleneck will remain at WC1 because it is still taking 2.2 hours for one job although it is working at = 110% efficiency.